Q3 : X is the element has 4 isotopes . Their masses are 53 . 93888 g / mol , 52 . 94065g mol , 49 . 94605 g / mol , and 51 . 94051 g / mol . X and X have a combined relative abundance of 87 . 9 % , and the X isotope has an relative abundance of 2 . 4 % . Given that the average atomic mass of X is 51 . 9961 g / mol a . Determine to the abundance of the Second isotope ? b . Determine to the abundance of the last 2 isotopes .
1
Expert's answer
2020-04-07T13:47:58-0400

a. x - abundance of the second isotope

(0.879 - x) * 53.93888 + x * 52.94065 + 0.024 * 49.94605 + (1 - 0.879 - 0.024) * 51.94051 = 51.9961

0.99823 * x = 1.65311

x = 1.656 not possible

(0.879 - x) * 51.94051 + x * 52.94065 + 0.024 * 53.93888 + 0.097 * 49.94065 = 51.9961

1.00015 * x = 0.2016

x = 0.20157 or 20.15%

b. 87.9 - 20.15 = 67.75%

67.75 + 9.7 = 77.45%